# TextRegion: Text-Aligned Region Tokens from Frozen Image-Text Models

### v1

@@ 1 Introduction
Contrastive image-text models, such as CLIP [25], enable open vocabulary retrieval and classification for images. However, these models typically operate at the global image level. What if we are interested in classifying, localizing, or retrieving specific regions given natural language text? The ability to understand precisely at the region level remains critical for many practical vision-language tasks, including open-vocabulary semantic segmentation, visual grounding, and fine-grained retrieval.
To achieve such detailed understanding, existing methods typically fall into two categories. One group of approaches, represented by Grounding DINO [20] and DenseCLIP [26], explicitly trains models using detailed annotations (e.g., bounding boxes or pixel-level masks). While effective, these training-based approaches struggle to match the broad concept coverage provided by large-scale, contrastively trained image-text models, making their lack of open-vocabulary object recognition a significant limitation for real-world applications. In contrast, methods in the second category, such as MaskCLIP [49], explore how to use existing components of the network to directly empower image-level CLIP for object-text alignment, without applying any training. However, patch-level embeddings are insufficient for retrieval or grounding tasks [30], as they often lack spatial coherence and clear boundary alignment.
Our TextRegion approach provides region-text alignment using the attention mechanism to aggregate patch features within each SAM [11] segmented region, similar to how the image-level CLS token is computed for image-text alignment. We also investigate the use of multi-resolution features and propose a method to mitigate the effect of “global” tokens that are prevalent in large models. Together, we show that our approach enables CLIP [25], SigLIP2 [35], and Perception Encoder [1] models to generate region-level tokens that are comparable to text embeddings, often outperforming more complex methods that require additional models or retraining.
@@ 2 Related Work
Region-Level Representation. Although patch-based representations dominate current vision research following the advent of vision transformers [36], region-level representations offer greater semantic richness and sparsity, often leading to superior performance in tasks such as segmentation, retrieval, and classification [30]. Recognizing their suitability for spatially detailed tasks, recent efforts have used region tokens for vision tasks [4, 30, 24, 15, 34]. However, existing methods typically involve task-specific training. Training-free region-based methods primarily targeted tasks that do not require language understanding, such as image retrieval [12, 32]. To the best of our knowledge, ours is the first study to explore generating text-aligned region tokens without training.
Open-Vocabulary Segmentation. Recent advances in open-world segmentation like MaskCLIP [49], SCLIP [38], and CLIPtrase [28] use image-text models to achieve zero-shot segmentation by aligning patch embeddings with textual descriptions. ProxyCLIP [14] enhances this by incorporating semantic-rich features from models like DINO [3] and DINOv2 [23]. The current SoTA, Trident [29] further integrates segmentation architectures like SAM [11] and SAM2 [27]. However, combining all three large models makes their framework complex. Unlike these patch-level methods, our approach aggregates patch embeddings into region-level features, effectively transforming the dense segmentation task into a sparse region classification problem.
Referring Expression Comprehension. This task involves identifying image regions described by textual queries. Recent zero-shot methods [33, 43, 42, 31] combine proposal boxes from MAttNet [45] with image-text models for image-caption matching. Evaluating spatial relation understanding is another important aspect of this task, leading some studies to develop methods that enhance spatial reasoning in image-text models [33]. However, our primary goal is to assess the visual-language alignment capability of our region tokens; thus, we focus exclusively on semantic retrieval without employing techniques to recover spatial relations.
Image-Text Models. Contrastive image-text embedding models, such as CLIP [25], ALIGN [9], SigLIP [46], SigLIP2 [35], and Perception Encoder [1], align visual and textual representations through dual encoders. These models generate text-aligned image embeddings, facilitating effective zero-shot performance across classification, text-to-image, and image-to-text retrieval. In this work, we investigate how their image-level embeddings can be transformed into region-level tokens, thus extending their applicability from global image understanding to precise region recognition.
@@ 3.1 Background and Notation
Contrastive image-text models, such as CLIP [25], learn to produce an image-level token [CLS] that is comparable to an embedding of text that captions the image.
The model divides the input image into $N$ patches and uses a convolutional layer to map them to patch embeddings, $[{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}}]\in\mathbb{R}^{N\times d}$, where $d$ is the embedding dimension. After adding positional embeddings, a learnable class token $\mathbf{x}_{\text{CLS}}\in\mathbb{R}^{d}$ is prepended to the sequence, resulting in $\mathbf{X}=[\mathbf{x}_{\text{CLS}};\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{(N+1)\times d}$. This sequence is processed through $L$ transformer layers. At each layer $l\in[1,L]$, the class token $\mathbf{x}_{\text{CLS}}$ progressively aggregates global context via self-attention. The attention projections are computed as, $\mathbf{Q}=\mathbf{X}\mathbf{W}_{q},\quad\mathbf{K}=\mathbf{X}\mathbf{W}_{k},\quad\mathbf{V}=\mathbf{X}\mathbf{W}_{v}$ where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ denote the learned projection matrices for queries, keys, and values. The attention weights are represented by $\boldsymbol{\alpha}\in\mathbb{R}^{(N+1)\times(N+1)}$. The attention output of the class token, denoted $\mathbf{y}_{\text{CLS}}$, is computed as a weighted sum over all value vectors $v$, where the weights $\alpha_{\text{cls},i}$ correspond to the attention scores between $\mathbf{x}_{\text{CLS}}$ and each patch token $\mathbf{x}_{i}$.
$\boldsymbol{\alpha}=\text{Softmax}\left(\frac{\mathbf{Q}(\mathbf{K})^{\top}}{\sqrt{d}}\right),\quad\mathbf{y}_{\text{CLS}}=\alpha_{\text{cls},\text{cls}}v_{\text{cls}}+\sum_{i=1}^{N}\alpha_{\text{cls},i}v_{i}=\boldsymbol{\alpha}_{\text{cls}}\mathbf{V},\quad\mathbf{y}_{\text{CLS}}\in\mathbb{R}^{d}$ (1)
In summary, the final-layer attention output $\mathbf{y}_{\text{CLS}}$ is a weighted sum over the value vectors of all patch tokens, with attention weights $\boldsymbol{\alpha}_{\text{cls}}$ determining each patch’s contribution to the image [CLS] token.
@@ 3.2 TextRegion Approach
Figure 3: TextRegion Framework. Mask Generation: We generate $R$ soft masks using SAM2, with values ranging from 0 to 1, where each mask corresponds to a distinct region in the input image. Patch Encoding: The image is encoded to obtain a multi-resolution feature map, which is fed into the final attention block of the frozen image-text models. See Sec. 4 for details. Mask-based Attention Pooling: As illustrated in Fig. 2, we perform pooling based on the $R$ bilinearly downsampled masks. Prediction: Using the pooled text-aligned region tokens, we support both zero-shot region-sparse classification and dense prediction.
Key Insight. Echoing MaskCLIP [49], our analysis of Eq. 1 suggest that the projected values from the value projection layer in the final block are quite rich in linguistic semantics, whereas the attention weights $\boldsymbol{\alpha}_{\text{cls}}$ is trained to capture global information rather than local information. To enhance the region-text alignment, one key is how to organize the values. To this end, we impose region-specific attention constraints, restricting the [CLS] token to attend only to patches within a designated region of interest. We present the details of region-specific attention constraints in Fig. 2. Basically, this involves ignoring patches that are unrelated to the target region. Since most patch values within a region share the same segmentation prediction, simply pooling these patch values to form a region token can enhance the robustness of the visual tokens.
@@
Mask Generation. For an input image of size $H\times W$, we apply SAM2 [27] to segment it into $R$ regions. This yields a set of soft masks $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{R}\right]\in\mathbb{R}^{H\times W}$, where each mask $\mathbf{M}_{r}$ contains per-pixel logits $M_{h,w}\in[0,1]$. A higher value of $M_{h,w}$ indicates that pixel $(h,w)$ is more likely to belong to region $r$. We downsample the high-resolution soft mask $\mathbf{M}_{r}$ to the feature map resolution of the vision encoder using bilinear interpolation, obtaining $\mathbf{m_{r}}\in\mathbb{R}^{N}$. $N$ means the number of patch tokens. Each element $m_{r,i}\in[0,1]$ represents the relevance of patch $i$ to region $r$.
@@
Some models, such as SigLIP [46], compute a delegate [CLS] token using a final attention pooling block. Our approach also applies to these models with slight implementation differences, which we detail in the supplemental material.
Prediction. With the global image zero-shot classification ability maintained, TextRegion supports both zero-shot region classification and pixel-level prediction. For region classification, we encode candidate text labels into embeddings $\mathbf{e}_{\text{text}}$ using the pretrained text encoder. Region classification logits are computed by cosine similarity between the region token $\mathbf{y}_{\text{r}}$ and each text embedding, weighted by temperature $\gamma$ ($=100$ in CLIP):
@@ 3.3 Region Token Refinement
Reducing global token interference. As illustrated in Fig. 4, we find that the large model often successfully segments more challenging images but fails on simpler cases. We attribute this to the presence of global patches [6, 28], whose tokens store global image information rather than patch-level local information. When aggregated via mask-based attention, such patches bias region tokens toward global text semantics, degrading localization performance. We introduce region consistency validation to filter out global patches. For each patch $i$ within region $r$ (as defined by the mask $\mathbf{M}_{r}$), we compute both intra-region and inter-region similarity scores:
@@
Here, $\mathbf{P_{r}}$ and $\mathbf{P}_{\neg r}$ represent the sets of patches inside and outside region $r$, respectively. A patch is identified as global if its local similarity $S_{\text{local }}=s_{\mathrm{in},i}-s_{\mathrm{out},i}<\tau$, where $\tau$ is a predefined threshold. Such patches are excluded from region token aggregation, ensuring that region-specific attention remains focused on locally relevant patch tokens. We apply a fixed threshold $\tau=0.07$ across all images, but exploring a dynamic thresholding strategy could be a valuable future work.
Multi-resolution patch encoding. Standard image-text models process fixed-resolution inputs, producing low-dimensional value feature maps that poorly capture small regions, often resulting in ineffective region masks (near-zero activations). Inspired by LLaVA’s AnyResolution [19] strategy, we propose: (1) split the original image into non-overlapping crops, resize and encode each to obtain localized features, and then concatenate them into a high-resolution feature map $V_{\text{high}}$; (2) concurrently, encode the full image to retain global context but low-resolution feature $V_{\text{low}}$. The final feature $V_{\text{final}}=V_{\text{high}}+\text{upsample}\left(V_{\text{low}}\right)$, which integrates local details with global semantics context, improving representation of both small and large regions.
@@ 4 Experiments
We show the strong region classification capabilities of TextRegion in Sec. 4.1, where our text-aligned region tokens outperform more complex methods that rely on additional models or retraining. In Sec. 4.2, we show that TextRegion also effectively handles instance queries, i.e., referring expression comprehension. Moreover, our framework seamlessly integrates with other image-text models, such as SigLIP2 and Perception Encoder, to produce meaningful text-aligned region tokens. Going beyond standard referring expression comprehension, which retrieves only a single region, we further demonstrate in Sec. 4.3 that our pipeline supports multiple object grounding. For all experiments, we use SAM2 [27] with the Hiera-Large backbone to generate region masks. Additional hyperparameters are detailed in the supplemental material.
@@ 4.1 Open-world Semantic Segmentation
Dataset. We evaluate on six widely used semantic segmentation benchmarks: PASCAL VOC 2012 [8], PASCAL Context [22], COCO-Stuff [2], COCO-Object [18], Cityscapes [5], ADE20K [48]. For VOC and Context, we test the settings without (VOC20, Context59) and with (Voc21, Context60) background labels. Since the image resolutions vary across datasets, we follow the design of previous works by setting different resizing scales for each dataset. Specifically, the shorter side is resized to 672 pixels for VOC20, VOC21, COCO-Object, Context59, ADE, and Context60; 896 pixels for COCO-Stuff; and 1344 pixels for Cityscapes.
@@
DINO [3] SAM VOC21 Context60 Object VOC20 Context59 Stuff City ADE
@@
TTD [10] 61.1 37.4 37.4 - - 23.7 27.0 17.0 -
CLIP-DINOiser [40] 62.1 32.4 34.8 80.9 35.9 24.6 31.7 20.0 40.3
SAM-CLIP [39] 60.6 29.2 - - - 31.5 - 17.1 -
@@
MaskCLIP [7] 43.4 23.3 20.6 74.9 26.4 16.7 24.9 11.9 30.3
CLIP Surgery [16] - - - - - 21.9 31.4 - -
SCLIP [37] 59.1 30.4 30.5 80.4 34.2 22.4 32.2 16.1 38.2
ProxyCLIP [14] 61.3 35.3 37.5 80.3 39.1 26.5 38.1 20.2 42.3
Trident [29] 67.1 38.6 41.1 84.5 42.2 28.3 42.9 21.9 45.8
@@
MaskCLIP [7] 31.4 13.3 16.2 41.7 15.8 8.4 17.7 10.4 19.3
SCLIP [37] 43.8 23.5 24.6 67.5 25.6 16.8 19.5 11.3 29.1
ProxyCLIP [14] 65.0 35.4 38.6 83.3 39.6 26.8 42.0 24.2 44.4
Trident [29] 70.8 40.1 42.2 88.7 44.3 28.6 47.6 26.7 48.6

### v2

@@ 1 Introduction
Contrastive image-text models, such as CLIP (radford2021learning), enable open vocabulary retrieval and classification for images. However, these models typically operate at the global image level. What if we are interested in classifying, localizing, or retrieving specific regions given natural language text? The ability to understand precisely at the region level remains critical for many practical vision-language tasks, including open-vocabulary semantic segmentation, visual grounding, and fine-grained retrieval.
To achieve such detailed understanding, existing methods typically fall into two categories. One group of approaches, represented by Grounding DINO (liu2024grounding) and DenseCLIP (rao2022denseclip), explicitly trains models using detailed annotations (e.g., bounding boxes or pixel-level masks). While effective, these training-based approaches struggle to match the broad concept coverage provided by large-scale, contrastively trained image-text models, making their lack of open-vocabulary object recognition a significant limitation for real-world applications. In contrast, methods in the second category, such as MaskCLIP (zhou2022extract), explore how to use existing components of the network to directly empower image-level CLIP for object-text alignment, without applying any training. Those works show that although image-text models are trained primarily to understand images at a global level, they implicitly acquire some ability for fine-grained visual–textual matching. However, patch-level embeddings are insufficient for retrieval or grounding tasks (Shlapentokh-Rothman_2024_CVPR), as they often lack spatial coherence and clear boundary alignment.
Given that modern segmentation models such as SAM (kirillov2023segment) provide precise spatial boundaries, while image-text models like CLIP (radford2021learning) capture what objects are present in images, an important question arises: can we combine the strengths of these two families of foundation models to achieve detailed region-level understanding? Moreover, as most prior work on detailed understanding has concentrated on CLIP, it is still uncertain whether alternative image-text encoders — such as SigLIP (zhai2023sigmoid) can effectively address this task.
In this work, we show that when supplied with accurate object masks to resolve spatial boundaries, image–text models can achieve strong zero-shot performance on detailed understanding tasks. Our TextRegion approach provides region-text alignment using the attention mechanism to aggregate patch features within each SAM segmented region, similar to how the image-level CLS token is computed for image-text alignment. We also investigate the use of multi-resolution features and propose a method to mitigate the effect of “global” tokens that are prevalent in large models. Together, TextRegion enables CLIP (radford2021learning), SigLIP2 (tschannen2025siglip), and Perception Encoder (bolya2025perception) models to generate region-level tokens that are comparable to text embeddings, often outperforming more complex methods that require additional models or retraining.
@@ 2 Related Work
Region-Level Representation. Although patch-based representations dominate current vision research following the advent of vision transformers (vaswani2017attention), region-level representations offer greater semantic richness and sparsity, often leading to superior performance in tasks such as segmentation, retrieval, and classification (Shlapentokh-Rothman_2024_CVPR). Recognizing their suitability for spatially detailed tasks, recent efforts have used region tokens for vision tasks (cheng2024spatialrgpt; Shlapentokh-Rothman_2024_CVPR; pan2024tokenize; lee2024toward; sun2024alpha). However, existing methods typically involve task-specific training. Training-free region-based methods primarily targeted tasks that do not require language understanding, such as image retrieval (korfhage2024search; sidhu2024search). To the best of our knowledge, ours is the first study to explore generating text-aligned region tokens without training.
Open-Vocabulary Segmentation. Recent advances in open-world segmentation like MaskCLIP (zhou2022extract), SCLIP (wang2025sclip), ClearCLIP (wang2025sclip) and CLIPtrase (lan2024clearclip) use image-text models to achieve zero-shot segmentation by aligning patch embeddings with textual descriptions. ProxyCLIP (lan2024proxyclip) enhances this by incorporating semantic-rich features from models like DINO (caron2021emerging) and DINOv2 (oquab2023dinov2). The current SoTA, Trident (shi2024harnessing) further integrates segmentation architectures like SAM (kirillov2023segment) and SAM2 (ravi2024sam). However, combining all three large models makes their framework complex. Unlike these patch-level methods, our approach aggregates patch embeddings into region-level features, effectively transforming the dense segmentation task into a sparse region classification problem.
Referring Expression Comprehension. This task involves identifying image regions described by textual queries. Recent zero-shot methods (subramanian2022reclip; yao2024cpt; yang2023fine; shtedritski2023does) combine proposal boxes from MAttNet (yu2018mattnet) with image-text models for image-caption matching. Evaluating spatial relation understanding is another important aspect of this task, leading some studies to develop methods that enhance spatial reasoning in image-text models (subramanian2022reclip). However, our primary goal is to assess the visual-language alignment capability of our region tokens; thus, we focus exclusively on semantic retrieval without employing techniques to recover spatial relations.
Image-Text Models. Contrastive image-text embedding models, such as CLIP (radford2021learning), ALIGN (jia2021scaling), SigLIP (zhai2023sigmoid), SigLIP2 (tschannen2025siglip), and Perception Encoder (bolya2025perception), align visual and textual representations through dual encoders. These models generate text-aligned image embeddings, facilitating effective zero-shot performance across classification, text-to-image, and image-to-text retrieval. In this work, we investigate how their image-level embeddings can be transformed into region-level tokens, thus extending their applicability from global image understanding to precise region recognition.
@@ 3.1 Background and Notation
Contrastive image-text models, such as CLIP (radford2021learning), learn to produce an image-level token [CLS] that is comparable to an embedding of text that captions the image.
The model divides the input image into $N$ patches and uses a convolutional layer to map them to patch embeddings, ${\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{N}}]\in\mathbb{R}^{N\times d}$, where $d$ is the embedding dimension. After adding positional embeddings, a learnable class token $\mathbf{x}_{\text{CLS}}\in\mathbb{R}^{d}$ is prepended to the sequence, resulting in $\mathbf{X}=[\mathbf{x}_{\text{CLS}};\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{N}]\in\mathbb{R}^{(N+1)\times d}$. This sequence is processed through $L$ transformer layers. At each layer $l\in[1,L]$, the class token $\mathbf{x}_{\text{CLS}}$ progressively aggregates global context via self-attention. The attention projections are computed as, $\mathbf{Q}=\mathbf{X}\mathbf{W}_{q},\quad\mathbf{K}=\mathbf{X}\mathbf{W}_{k},\quad\mathbf{V}=\mathbf{X}\mathbf{W}_{v}$ where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ denote the learned projection matrices for queries, keys, and values. The attention weights are represented by $\bm{\alpha}\in\mathbb{R}^{(N+1)\times(N+1)}$. The attention output of the class token, denoted $\mathbf{y}_{\text{CLS}}$, is computed as a weighted sum over all value vectors $v$, where the weights $\alpha_{\text{cls},i}$ correspond to the attention scores between $\mathbf{x}_{\text{CLS}}$ and each patch token $\mathbf{x}_{i}$.
$\bm{\alpha}=\text{Softmax}\left(\frac{\mathbf{Q}(\mathbf{K})^{\top}}{\sqrt{d}}\right),\quad\mathbf{y}_{\text{CLS}}=\alpha_{\text{cls},\text{cls}}v_{\text{cls}}+\sum_{i=1}^{N}\alpha_{\text{cls},i}v_{i}=\bm{\alpha}_{\text{cls}}\mathbf{V},\quad\mathbf{y}_{\text{CLS}}\in\mathbb{R}^{d}$ (1)
In summary, the final-layer attention output $\mathbf{y}_{\text{CLS}}$ is a weighted sum over the value vectors of all patch tokens, with attention weights $\bm{\alpha}_{\text{cls}}$ determining each patch’s contribution to the image [CLS] token.
@@
Figure 3: TextRegion Framework. Mask Generation: We generate $R$ soft masks using SAM2, with values ranging from 0 to 1, where each mask corresponds to a distinct region in the input image. Patch Encoding: The image is encoded to obtain a multi-resolution feature map, which is fed into the final attention block of the frozen image-text models. See Sec. 3.3 for details. Mask-based Attention Pooling: As illustrated in Fig. 2, we perform pooling based on the $R$ bilinearly downsampled masks. Prediction: Using the pooled text-aligned region tokens, we support both zero-shot region-sparse classification and dense prediction.
@@ 3.2 TextRegion Approach
Key Insight. Echoing MaskCLIP (zhou2022extract), Eq. 1 suggest that the projected values from the value projection layer in the final block are quite rich in linguistic semantics, whereas the attention weights $\bm{\alpha}_{\text{cls}}$ is trained to capture global information rather than local information. To enhance the region-text alignment, one key is how to organize the values. To this end, we impose region-specific attention constraints, restricting the [CLS] token to attend only to patches within a designated region of interest. We present the details of region-specific attention constraints in Fig. 2. Basically, this involves ignoring patches that are unrelated to the target region. Since most patch values within a region share the same segmentation prediction, simply pooling these patch values to form a region token can enhance the robustness of the visual tokens.
@@
Mask Generation. For an input image of size $H\times W$, we apply SAM2 (ravi2024sam) to segment it into $R$ regions. This yields a set of soft masks $\mathcal{M}=\left[\mathbf{M}_{1};\mathbf{M}_{2};\ldots;\mathbf{M}_{R}\right]\in\mathbb{R}^{R\times H\times W}$, where each mask $\mathbf{M}_{r}$ contains per-pixel logits $M_{h,w}\in[0,1]$. A higher value of $M_{h,w}$ indicates that pixel $(h,w)$ is more likely to belong to region $r$. We downsample the high-resolution soft mask $\mathbf{M}_{r}$ to the feature map resolution of the vision encoder using bilinear interpolation, obtaining $\mathbf{m_{r}}\in\mathbb{R}^{N}$. $N$ means the number of patch tokens. Each element $m_{r,i}\in[0,1]$ represents the relevance of patch $i$ to region $r$.
@@
Some models, such as SigLIP (zhai2023sigmoid), compute a delegate [CLS] token using a final attention pooling block. Our approach also applies to these models with slight implementation differences, which we detail in Appendix Sec A.3.
Prediction. With the global image zero-shot classification ability maintained, TextRegion supports both zero-shot region classification and pixel-level prediction. For region classification, we encode candidate text labels into embeddings $\mathbf{e}_{\text{text}}$ using the pretrained text encoder. Region classification logits are computed by cosine similarity between the region token $\mathbf{y}_{\text{r}}$ and each text embedding, weighted by temperature $\gamma$ (=100=100 in CLIP):
@@ 3.3 Region Token Refinement
Reducing global token interference. As illustrated in Fig. 4, we find that the large model often successfully segments more challenging images but fails on simpler cases. We attribute this to the presence of global patches. shao2024explore; darcet2023vision and yang2025visionzip observe that certain tokens in vision transformers capture global image semantics rather than localized patch-level details. When aggregated via mask-based attention, such patches bias region tokens toward global text semantics, degrading localization performance. We introduce region consistency validation to filter out global patches. For each patch $i$ within region $r$ (as defined by the mask $\mathbf{M}_{r}$), we compute both intra-region and inter-region similarity scores:
@@
Here, $\mathbf{P_{r}}$ and $\mathbf{P}_{\neg r}$ represent the sets of patches inside and outside region $r$, respectively. A patch is identified as global if its local similarity $S_{\text{local }}=s_{\mathrm{in},i}-s_{\mathrm{out},i}<\tau$, where $\tau$ is a predefined threshold. Such patches are excluded from region token aggregation, ensuring that region-specific attention remains focused on locally relevant patch tokens. We apply a fixed threshold $\tau=0.07$ across all images.
Multi-resolution patch encoding. Standard image-text models process fixed-resolution inputs, producing low-dimensional value feature maps that poorly capture small regions, often resulting in ineffective region masks (near-zero activations). Inspired by LLaVA’s AnyResolution (liu2024improved) strategy, we propose: (1) split the original image into non-overlapping crops, resize and encode each to obtain localized features, and then concatenate them into a high-resolution feature map $V_{\text{high}}$; (2) concurrently, encode the full image to retain global context but low-resolution feature $V_{\text{low}}$. The final feature $V_{\text{final}}=V_{\text{high}}+\text{upsample}\left(V_{\text{low}}\right)$, which integrates local details with global semantics context, improving representation of both small and large regions.
@@ 4 Experiments
We show the strong region classification capabilities of TextRegion in Sec. 4.1, where our text-aligned region tokens outperform more complex methods that rely on additional models or retraining. In Sec. 4.2, we show that TextRegion also effectively handles instance queries, i.e., referring expression comprehension. Moreover, our framework seamlessly integrates with other image-text models, such as SigLIP2 and Perception Encoder, to produce meaningful text-aligned region tokens. Going beyond standard referring expression comprehension, which retrieves only a single region, we further demonstrate in Sec. 4.3 that our pipeline supports multiple object grounding. For all experiments, we use SAM2 (ravi2024sam) with the Hiera-Large backbone to generate region masks. Additional hyperparameters are detailed in Appendix Sec A.2.
@@ 4.1 Open-world Semantic Segmentation
Dataset. We evaluate on six widely used semantic segmentation benchmarks: PASCAL VOC 2012 (everingham2015pascal), PASCAL Context (mottaghi2014role), COCO-Stuff (caesar2018coco), COCO-Object (lin2014microsoft), Cityscapes (cordts2016cityscapes), ADE20K (zhou2019semantic). For VOC and Context, we test the settings without (VOC20, Context59) and with (Voc21, Context60) background labels. Since the image resolutions vary across datasets, we follow the design of previous works by setting different resizing scales for each dataset. Specifically, the shorter side is resized to 672 pixels for VOC20, VOC21, COCO-Object, Context59, ADE, and Context60; 896 pixels for COCO-Stuff; and 1344 pixels for Cityscapes.
@@
DINO SAM VOC21 Context60 Object VOC20 Context59 Stuff City ADE
@@
TTD (jo2024ttd) 61.1 37.4 37.4 - - 23.7 27.0 17.0 -
CLIP-DINOiser (wysoczanska2024clip) 62.1 32.4 34.8 80.9 35.9 24.6 31.7 20.0 40.3
SAM-CLIP (wang2024sam) 60.6 29.2 - - - 31.5 - 17.1 -
@@
MaskCLIP (dong2023maskclip) 43.4 23.3 20.6 74.9 26.4 16.7 24.9 11.9 30.3
SCLIP (wang2024sclip) 59.1 30.4 30.5 80.4 34.2 22.4 32.2 16.1 38.2
ResCLIP (yang2025resclip) - - - 86.0 36.8 24.7 35.9 18.0 -
ProxyCLIP (lan2024proxyclip) 61.3 35.3 37.5 80.3 39.1 26.5 38.1 20.2 42.3
LaVG (kang2024defense) 62.1 31.6 34.2 82.5 34.7 23.2 25.0 15.8 38.6
LPOSS+ (stojnic2025lposs) 62.4 35.4 34.3 79.3 38.6 26.5 37.9 22.3 42.1
Trident (shi2024harnessing) 67.1 38.6 41.1 84.5 42.2 28.3 42.9 21.9 45.8
CLIPtrase (shao2024explore) 57.1 32.0 44.2 82.2 36.4 24.8 - - -
@@
SCLIP (wang2024sclip) 43.8 23.5 24.6 67.5 25.6 16.8 19.5 11.3 29.1
ProxyCLIP (lan2024proxyclip) 65.0 35.4 38.6 83.3 39.6 26.8 42.0 24.2 44.4
Trident (shi2024harnessing) 70.8 40.1 42.2 88.7 44.3 28.6 47.6 26.7 48.6